Consider two alternatives: B A $300 $500 $75 Cost $75 Uniform annual benefit X Infinity Useful life, in years Assume that AltPlease don't use excel and show all work.
Answer #1

AW, Project A ($) = - 500 x 0.1 + 75 = - 50 + 75 = 25

AW, Project B = - 300 x A/P(10%, X) + 75 = 25

- 300 x A/P(10%, X) = - 50

A/P(10%, X) = 1/6

0.1 / [1 - (1.1)-X]** = 1/6

[1 - (1.1)-X] = 0.6

(1.1)-X = 0.4

(1.1)X = 2.5

Taking natural logarithm on each side,

X x ln 1.1 = ln 2.5

X x 0.0953 = 0.9163

X = 9.61 years

**A/P(r%, N) = r / [1 - (1 + r)-N]